# Symmetry and Topology in the Non-Hermitian Kitaev chain

### v1

@@ IV Conclusion
The analysis presented here focuses on the non-Hermitian Kitaev chain where non-Hermiticity is introduced via non-reciprocal hopping amplitudes ($t_{1}\neq t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}\neq\Delta_{2}e^{i\varphi_{2}}$), cf. Eq. (III). Considering arbitrary complex parameters, we show the existence of different discrete symmetries corresponding to each parameter regime, see Tab. 1.
We demonstrated that the transition point between the topological and trivial phase with periodic boundary condition (equivalently infinite chain) coincides with the open boundary condition for a finite chain. The underlying reason is the existence of particle-hole symmetry across entire parameter regime, see Tab. 1. As a result, the solutions to the localization parameter $\beta$, Eq. (9) for the zero modes (doubly degenerate) of the Hamiltonian always exist in reciprocal pairs Li_2023; Yan_2023; https://doi.org/10.1002/pssb.202200448 and therefore zero-modes are robust to non-Hermitian skin effect (NHSE), see Fig. 3a-c and Fig. 4a. We also study the bulk (non-zero energy) modes and find that the particle and hole wavefunctions show equal and opposite skin effect, as implied by PHS, see Fig. 4b.
@@
With recent experimental realization of non-Hermitian skin effect in mechanical metamaterials doi:10.1073/pnas.2010580117; PhysRevLett.131.207201, topoelectric circuits zou_observation_2021; 7dvt-k6gk and photonic systems doi:10.1126/science.aaz8727; Xiao2020; PhysRevResearch.2.013280; zhang_acoustic_2021; gao_non-hermitian_2021; zhou_observation_2023; wan_observation_2023, the interplay between topology and non-Hermiticity has become experimentally relevant. The versatility of these platforms allows for the engineering of a wide range of non-reciprocal couplings and effective interactions https://doi.org/10.1002/lpor.202400099, making the realization of non-Hermitian Majorana physics a tangible prospect. In this work, we have provided a comprehensive characterization of the non-Hermitian Kitaev chain over the full complex parameter space, establishing a consistent correspondence between bulk topology, zero-energy Majorana modes, and phase boundaries. For specific parameter regimes, we obtained analytical expressions for points on the topological phase boundary, as illustrated in Fig. 5, offering concrete insight into the structure of the phase diagram. It remains to be seen whether the phase diagram can be determined analytically in full generality, including a characterization of the shape of the phase boundary; nevertheless, our results lay a systematic and broadly applicable foundation for exploring non-Hermitian topological phase transitions in this and related models.
@@ Acknowledgements
We thank Jukka I. Vayrynen and Qi Zhou for insightful comments and discussions. A.R. would also like to thank Laimei Nie and Eric Schultz for useful discussions.
@@ Appendix A Symmetry analysis for non-Hermitian Kitaev chain
and $\mathcal{C}_{k}^{\dagger}=\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}$. The discrete symmetries have been discussed in Kawabata_2022. The spectral constraints arising from the possible symmetries of the system are summarized in the Tab. A1.
@@
We now explore the possible symmteries of the non-Hermitian Kitaev chain and obtain Tab. 1 in the main text. The symmetries are unitary relations between $2\times 2$ matrices, and a necessary and sufficient check is given by Specht’s theorem, which states that two $2\times 2$ matrices $A$ and $B$ are unitarily equivalent if and only if the following 3 criteria are satisfied

### v2

@@ IV Conclusion
The analysis presented here focuses on the non-Hermitian Kitaev chain where non-Hermiticity is introduced via non-reciprocal hopping amplitudes ($t_{1}\neq t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}\neq\Delta_{2}e^{i\varphi_{2}}$), cf. Eq. (III). Considering arbitrary complex parameters, we show the existence of different discrete symmetries corresponding to each parameter regime, see Table 1.
We demonstrated that the transition point between the topological and trivial phase with periodic boundary condition (equivalently infinite chain) coincides with the open boundary condition for a finite chain. The underlying reason is the existence of particle-hole symmetry across entire parameter regime, see Table 1. As a result, the solutions to the localization parameter $\beta$, Eq. (9) for the zero modes (doubly degenerate) of the Hamiltonian always exist in reciprocal pairs [41, 69, 59] and therefore zero-modes are robust to non-Hermitian skin effect (NHSE), see Fig. 3(a)-(c) and Fig. 4(a). We also study the bulk (non-zero energy) modes and find that the particle and hole wavefunctions show equal and opposite skin effect, as implied by PHS, see Fig. 4(b).
@@
With recent experimental realization of non-Hermitian skin effect in mechanical metamaterials [22, 63], topoelectric circuits [81, 18] and photonic systems [64, 67, 80, 76, 21, 79, 62], the interplay between topology and non-Hermiticity has become experimentally relevant. The versatility of these platforms allows for the engineering of a wide range of non-reciprocal couplings and effective interactions [78], making the realization of non-Hermitian Majorana physics a tangible prospect. In this work, we have provided a comprehensive characterization of the non-Hermitian Kitaev chain over the full complex parameter space, establishing a consistent correspondence between bulk topology, zero-energy Majorana modes, and phase boundaries. For specific parameter regimes, we obtained analytical expressions for points on the topological phase boundary, as illustrated in Fig. 5, offering concrete insight into the structure of the phase diagram. It remains to be seen whether the phase diagram can be determined analytically in full generality, including a characterization of the shape of the phase boundary; nevertheless, our results lay a systematic and broadly applicable foundation for exploring non-Hermitian topological phase transitions in this and related models.
@@ Acknowledgements
We thank Jukka I. Vayrynen and Qi Zhou for insightful comments and discussions. A.R. would also like to thank Laimei Nie and Eric Schultz for useful discussions. We also thank Sayak Ray for comments on the draft of the manuscript.
@@ Appendix A Symmetry analysis for non-Hermitian Kitaev chain
and $\mathcal{C}_{k}^{\dagger}=\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}$. The discrete symmetries have been discussed in [33]. The spectral constraints arising from the possible symmetries of the system are summarized in the Table A1.
@@
We now explore the possible symmteries of the non-Hermitian Kitaev chain and obtain Tabel 1 in the main text. The symmetries are unitary relations between $2\times 2$ matrices, and a necessary and sufficient check is given by Specht’s theorem, which states that two $2\times 2$ matrices $A$ and $B$ are unitarily equivalent if and only if the following 3 criteria are satisfied